### v1

# Improving Requirements Classification with SMOTE-Tomek Preprocessing

Abstract
This study emphasizes the domain of requirements engineering by applying the SMOTE-Tomek preprocessing technique, combined with stratified K-fold cross-validation, to address class imbalance in the PROMISE dataset. This dataset comprises 969 categorized requirements, classified into functional and non-functional types. The proposed approach enhances the representation of minority classes while maintaining the integrity of validation folds, leading to a notable improvement in classification accuracy. Logistic regression achieved 76.16% ± 2.58%, significantly surpassing the baseline of 58.31% ± 2.05%. These results highlight the applicability and efficiency of machine learning models as scalable and interpretable solutions.
Index Terms:
Requirement Engineering, Class Imbalance, SMOTE-Tomek, Imbalanced Datasets, Requirement Classification, Natural Language Processing, Synthetic Oversampling, Text-Based Classification, Supervised Learning.
I Introduction
Requirements engineering (RE) is a cornerstone of the software development lifecycle, translating stakeholder needs into actionable system specifications [1, 2]. Accurate and efficient requirement classification is essential to ensure project clarity, prioritization, and goal alignment. Requirements are typically categorized into functional, non-functional, and subtypes [3, 4, 5, 6]. A survey revealed that over 60% of failed projects neglected non-functional requirements, underscoring the critical importance of effective classification [7]. Despite its significance, requirement classification remains a largely manual process, prone to inconsistencies, inefficiencies, and scalability challenges [8].
Over the years, researchers have explored various approaches to automate requirement classification. Early rule-based systems were among the first attempts, offering interpretable but inflexible solutions that were labor-intensive to maintain [9]. The advent of machine learning (ML) brought more adaptive models, demonstrating moderate success, particularly in structured datasets [10, 3, 6]. However, these methods often face challenges with class imbalance, a prevalent issue in real-world datasets, including the widely used PROMISE dataset for requirements engineering [11]. This imbalance skews model predictions toward majority classes, undermining performance on minority classes.
Over the past decade, deep learning (DL) has achieved transformative breakthroughs across multiple domains, including natural language processing (NLP), computer vision, and time series analysis. These advancements are primarily driven by the unprecedented representational power of deep neural networks, which leverage millions, and in some cases billions, of trainable parameters to extract and model intricate, high-dimensional patterns from large-scale datasets [12].
Equally significant are the advances in time series analysis, where DL models have demonstrated remarkable capabilities in tasks such as motion sensing classification and physical quantity estimation [13, 14, 15, 16]. These innovations highlight DL’s ability to capture temporal dependencies and complex properties in sequential data.
In computer vision, architectures such as ResNet [17] and Vision Transformers (ViT) [18] have revolutionized image classification, object detection, and semantic segmentation, achieving state-of-the-art accuracy while enabling applications that were previously unattainable.
In NLP, DL models such as BERT [19] and GPT-3 [20] have set new standards in tasks like language translation, text summarization, and question answering, surpassing traditional methods and enabling more nuanced understanding and generation of human language [21, 22, 23, 24, 25]. These models excel in NLP tasks due to their ability to capture contextual relationships within text using self-attention mechanisms. They are particularly effective at modeling semantic intricacies in textual data, enabling strong performance in complex classification tasks. However, these methods require not only substantial computational resources for fine-tuning and inference but also access to massive amounts of high-quality data to achieve their full potential. This reliance on extensive datasets and powerful hardware often makes them less practical for small-scale or resource-constrained projects, where data availability and computational capacity may be limited.
To address these challenges, this study integrates the Synthetic Minority Oversampling Technique (SMOTE) [26] with Tomek Links (SMOTE-Tomek) [27], providing a robust preprocessing solution for imbalanced datasets. SMOTE oversamples minority classes by generating synthetic samples through interpolation, effectively enhancing the diversity of underrepresented classes while mitigating overfitting, a common issue with random oversampling methods. Complementing this, Tomek Links identifies and removes noisy or borderline samples, improving class separability by reducing overlapping data points between classes. This dual-action preprocessing strategy results in cleaner and more balanced training datasets, enabling ML models to better learn decision boundaries for minority classes.
The proposed approach is particularly suited to text-based datasets like the PROMISE dataset, where class imbalance and noise frequently undermine classification performance. By addressing these issues, SMOTE-Tomek creates an optimal foundation for training modern ML models, which can achieve better performance without the computational demands of deep learning approaches. This study builds on the strengths of SMOTE-Tomek by integrating it into a stratified K-fold cross-validation framework, preserving the integrity of validation folds and ensuring rigorous, unbiased evaluation.
In RE classification, SMOTE-Tomek addresses both class imbalance and noise, making it particularly suitable for the PROMISE dataset, where minority classes are significantly underrepresented. While previous studies have demonstrated the efficacy of SMOTE-Tomek for addressing class imbalance in the PROMISE dataset [28], this research extends its application by integrating it into a stratified K-fold cross-validation framework. This ensures the validation folds remain untouched by resampling, preserving the original class distribution and enabling rigorous evaluation of the model’s generalization capabilities. Furthermore, this study explores several classical ML models, evaluates their performance on the PROMISE dataset, and highlights the potential of lightweight solutions for efficient requirement classification.
This study makes the following key contributions:
  • A systematic application of SMOTE-Tomek to address class imbalance in the PROMISE dataset, within a 10-fold cross-validation framework that ensures the validation set remains unaffected by resampling.
  • A comparative evaluation of classical ML models, demonstrating their potential for scalable and efficient requirement classification.
  • An analysis of Logistic Regression (LR) coefficients to interpret the relationship between features and requirement types, offering insights into the most frequent and influential words for each type.
The experimental results demonstrate that integrating SMOTE-Tomek with classical ML frameworks improves classification performance. LR, for instance, achieves an accuracy improvement from 58.31% ± 2.05% to 76.16% ± 2.58% under cross-validation.
The remainder of this paper is structured as follows: Section II outlines the methodology, including an overview of the dataset, preprocessing steps, the class imbalance challenge, the SMOTE-Tomek approach, and the classical ML models utilized. Section III presents the results and their implications, and Section IV concludes with directions for future research.
II Learning Method
This section provides an overview of the methodology employed in this study, detailing the dataset, class imbalance challenges, and the preprocessing techniques applied to address them. It outlines the evaluated ML models, the implementation of the SMOTE-Tomek approach for data balancing, and the use of K-fold cross-validation to ensure robust and unbiased performance assessment.
II-A Dataset
The study utilizes the expanded PROMISE dataset, a diverse collection of 969 software requirements extracted from Software Requirements Specification (SRS) documents using the Google search engine [11]. The dataset contains 444 functional requirements (45.8%) and 525 non-functional requirements (54.2%), distributed across 12 distinct categories. These categories represent a spectrum of requirements in software engineering, such as Security (SE), Usability (US), Portability (PO), and Performance (PE). Despite its comprehensiveness, the dataset exhibits notable class imbalance. For example, the Portability category is severely underrepresented, comprising only 12 requirements (1.24%), while the most prevalent non-functional class, Security, includes 125 requirements (12.9%). This disparity is further amplified in the broader classification between functional and non-functional requirements, which affects the performance and generalizability of ML models. Figure 1 illustrates the distribution of the dataset across all requirement types, highlighting the inherent imbalance.
Figure 1: Requirement type distribution in the PROMISE dataset.
The following is a concise description of the requirement types used in this study:
Functional Requirement (F)
Defines the specific functionalities and behaviors the system must perform to meet stakeholder needs.
Availability Requirement (A)
Ensures that the system remains operational and accessible under defined conditions and constraints.
Legal Requirement (L)
Addresses compliance with legal, regulatory, and contractual obligations relevant to the system’s deployment and use.
Look-and-Feel Requirement (LF)
Specifies the aesthetic and user interface characteristics to align with user expectations and branding.
Maintainability Requirement (MN)
Focuses on the system’s ability to be easily modified, updated, or repaired during its lifecycle.
Operability Requirement (O)
Ensures that the system can be operated efficiently and effectively within its intended environment.
Performance Requirement (PE)
Defines the system’s capability to deliver expected throughput, response time, and efficiency under specified conditions.
Scalability Requirement (SC)
Addresses the system’s ability to adapt to increased workloads or expanded functionalities without degradation.
Security Requirement (SE)
Protects the system and its data against unauthorized access, breaches, and vulnerabilities.
Usability Requirement (US)
Ensures the system is intuitive and user-friendly, enabling stakeholders to interact with it effectively.
Fault Tolerance Requirement (FT)
Ensures the system can continue functioning correctly even in the event of hardware or software failures.
Portability Requirement (PO)
Specifies the system’s ability to be deployed or operated across various platforms, environments, or configurations.
Table I provides examples of several requirements and their types from the PROMISE dataset.
TABLE I: Several Requirements and Their Types
From The PROMISE Dataset
Requirement Example Requirement Type
The product shall have audit capabilities. The product shall store messages for a minimum of one year for audit and transaction tracking purposes. Security (SE)
The recycled parts search results provided to the estimator shall be retrieved by the system. Functional (F)
Users should be able to access their streaming movies in under 2 clicks after logging into the website. Usability (US)
The product shall be available 24 hours per day, seven days per week. Availability (A)
The system shall help the user avoid making mistakes while scheduling classes and clinicals for the nursing students. Usability (US)
The system shall display a detailed invoice for the current order once it is confirmed. Functional (F)
The system will use a secured database. Security (SE)
The leads washing functionality will store any potential lead duplicates returned by the enterprise system. Functional (F)
The system shall interface with the faculty central server. Operability (O)
The product shall be internet browser independent. The product must run using Internet Explorer and Netscape Navigator. Maintainability (MN)
II-B Pre-Processing
Transforming unstructured text into a numerical representation suitable for ML algorithms is a fundamental preprocessing step in many NLP tasks. The PROMISE dataset comprises individual sentences, rather than full documents. The common Term Frequency-Inverse Document Frequency (TF-IDF) remains a highly effective vectorization technique [29, 30]. TF-IDF quantifies the importance of a term within a sentence while considering its prevalence across the entire dataset of sentences. By emphasizing terms that are frequent within a specific sentence but rare across the dataset, TF-IDF captures features that are contextually meaningful and relevant to the classification task. Despite the brevity of the textual units, TF-IDF preserves the semantic significance of terms.
II-C Class Imbalance Challenge
Class imbalance poses a critical challenge in training ML models, as they tend to prioritize majority classes, resulting in poor generalization and suboptimal performance on minority classes [31]. Such biases can significantly impact the usability and reliability of classification models in practical applications, where correctly identifying minority classes is often essential. Balancing the training set ensures that the model learns to give equal attention to all classes, preventing it from being overly biased toward majority classes.
II-D SMOTE-Tomek
Among various methods to address data imbalance, SMOTE-Tomek is a hybrid technique that combines the benefits of oversampling and data cleaning. The SMOTE addresses the imbalance by generating synthetic samples for minority classes, ensuring the training set has a more equitable class distribution. However, SMOTE alone can introduce noise by oversampling borderline or overlapping samples. Tomek Links complement SMOTE by identifying and removing borderline or ambiguous samples that may hinder model training. Together, SMOTE-Tomek enhances data quality while balancing the class distribution in the training phase. Unlike random oversampling, which duplicates existing minority class samples, SMOTE creates new synthetic samples based on linear interpolation between existing samples.
In practical applications of NLP tasks, where textual data is complex and often imbalanced, SMOTE-Tomek emerges as a robust and effective solution. The synthetic examples generated by SMOTE are not direct textual entities but instead constitute feature vectors that represent the underlying characteristics of the minority class, as presented in Table II. These vectors serve as abstract representations, enabling the model to internalize the structural properties of underrepresented classes without requiring additional annotated textual data.
The SMOTE-Tomek process is described in the following steps:
  1. 1.
    Generate Synthetic Samples for Minority Classes: For each sample $x_{i}$ belonging to the minority class, synthetic samples are generated using linear interpolation:
    $x_{\text{synthetic}}=x_{i}+\lambda(x_{nn}-x_{i}),$ (1)
    where $x_{nn}$ is a randomly selected nearest neighbor of $x_{i}$ from the same minority class, and $\lambda\in[0,1]$ is a random scalar.
  2. 2.
    Identify Tomek Links: A Tomek Link is defined as a pair of samples $x_{i}$ and $x_{j}thatare$ that are mutual nearest neighbors, meaning $x_{i}$ is the closest sample to $x_{j}$ and vice versa. Also, $x_{i}$ and $x_{j}$ are belonging to different classes, with one from the minority class and the other from the majority class. The equation for identifying Tomek links is give by:
    $d(x_{i},x_{j})=\min\{d(x_{i},x_{k})\mid x_{k}\},$ (2)
    where $d(x_{i},x_{j})$ is the distance (commonly Euclidean) between $x_{i}$ and $x_{j}$ and $x_{k}\in$ different class from $x_{i}$.
  3. 3.
    Remove Majority Class Samples: For each identified Tomek Link, remove the sample belonging to the majority class to enhance class separation and reduce noise.
TABLE II: Examples of Requirements and Types Resampling Using SMOTE-Tomek
Requirement Type
based, established, lead, operate, physical, process, product, run, server, service, shall, structure, washing, web O
changes, edit, information, just, modify, normal, personal, read, securely, server, transmission, transmitted, users SE
entry, shown F
initiator, mediator, messages, request, response, shall, user F
accommodate, compensatory, data, failures, fault, items, large, product, recovery, robust, routing, shall, technique, tolerance, transaction, using FT
explanatory, intuitive, self, shall US
control, details, employees, site, view F
browsers, ce, compatible, environments, expected, interface, major, operating, order, palm, product, run, standards, systems, usable, variety, web, wide, windows PO
800x600, 95, appropriately, correct, corrected, display, feel, higher, implementation, incorrect, intranet, look, notification, pages, prior, provide, remaining, resolutions, shall, uniform, web, week LF
able, alongside, days, environment, function, installed, java, operating, product, runtime, server, shall, software O
II-E Stratified K-fold Cross-Validation Method
Stratified K-fold cross-validation provides significant advantages over traditional train/test splits, particularly in handling class imbalance. By dividing the dataset into K equal folds and ensuring that each fold is used for validation exactly once, this method maximizes the use of available data. Through stratified sampling, it maintains the proportional representation of all classes in each fold, ensuring that minority classes are neither underrepresented nor overlooked in the training or validation sets.
To ensure a fair evaluation and maintain the integrity of the cross-validation process, SMOTE-Tomek resampling is applied solely to the training folds during each iteration of the 10-fold stratified cross-validation. Specifically, for each iteration, the resampling technique is applied to the 9 training folds to balance the classes, while the validation fold is left untouched by resampling. This preserves the true distribution of the original data in the validation fold, thereby avoiding data leakage and enabling a realistic evaluation of the model’s performance. The process, as illustrated in Figure 2, highlights that SMOTE-Tomek is applied only to the validation fold.
Figure 2: Cross validation with carefully applying SMOTE-Tomek only to the validation folds.
II-F Trainig Algorithm
The proposed method, as outlined in Algorithm 1, emphasizes the integration of stratified K-fold cross-validation to ensure class proportions are preserved across folds. Importantly, SMOTE-Tomek is applied solely to the training folds to prevent data leakage, thereby preserving the integrity of validation metrics. The algorithm also includes final training on the entire balanced dataset to fully utilize the available data. Table III presents the notations and their corresponding definitions as utilized in the algorithm.
TABLE III: Notations and Definitions
Notation Definition
$\mathcal{D}$ Complete dataset used for training and evaluation
$K$ Number of folds in stratified K-fold cross-validation
$\mathcal{F}_{i}$ Fold $i$, used as validation set in the $i$-th iteration
$\mathcal{D}_{\text{train}}$ Training set formed by $K-1$ folds
$\mathcal{D}_{\text{val}}$ Validation set corresponding to fold $i$
$\mathcal{M}_{i}$ Model trained in the $i$ iteration of cross-validation
Algorithm 1 Training with SMOTE-Tomek and Stratified K-Fold Cross-Validation
0:  $\mathcal{D}$, $K$, $\mathcal{M}$
1:  Split $\mathcal{D}$ into $K$ stratified folds $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$, ensuring class proportions are preserved
2:  for fold $i=1$ to $K$ do
3:     $\mathcal{D}_{\text{train}}\leftarrow\bigcup_{j\neq i}\mathcal{F}_{j}$Combine $K-1$ folds for training}
4:     $\mathcal{D}_{\text{val}}\leftarrow\mathcal{F}_{i}$Set $i$-th fold as the validation set}
5:     $\mathcal{D}_{\text{train\_balanced}}\leftarrow\text{Apply SMOTE-Tomek only to }\mathcal{D}_{\text{train}}$
6:     Train $\mathcal{M}_{i}$ on $\mathcal{D}_{\text{train\_balanced}}$
7:     Evaluate $\mathcal{M}_{i}$ on $\mathcal{D}_{\text{val}}$, compute metrics
8:  end for
9:  Compute average metrics across all $K$ folds
10:  Train $\mathcal{M}_{\text{final}}$ on $\mathcal{D}$ (balanced using SMOTE-Tomek)
11:  return  $\mathcal{M}_{\text{final}}$, average performance metrics
II-G Classical ML Models
The following ML models were evaluated in this study to capture diverse classification perspectives.
  • Decision Tree (DT): A tree-based algorithm that splits the data into subsets based on feature values, creating a hierarchy of decisions to classify data efficiently.
  • Random Forest: An ensemble method combining multiple DTs, where each tree votes, and the majority decision is taken for robust and accurate predictions.
  • Support Vector Machine (SVM) with Linear Kernel: A linear classifier that finds the hyperplane maximizing the margin between classes, optimized for linearly separable data.
  • SVM with RBF Kernel: A nonlinear classifier that maps data into a higher-dimensional space using the radial basis function (RBF) kernel for better class separation.
  • Naive Bayes: A probabilistic model based on Bayes’ theorem, assuming feature independence, and commonly used for text classification tasks.
  • Logistic Regression: A statistical model that predicts the probability of an outcome using a logistic function.
  • K-Nearest Neighbors (KNN): A non-parametric algorithm that classifies samples based on the majority class of their $k$-nearest neighbors in the feature space.
  • Gradient Boosting: An iterative ensemble method that builds sequential trees, correcting errors of previous trees to minimize classification error.
  • AdaBoost: An adaptive boosting algorithm that combines weak classifiers, assigning higher weights to misclassified samples to improve accuracy.
  • CatBoost: A gradient boosting algorithm optimized for categorical features.
III Results and Discussion
This section evaluates the performance of various ML models for requirement type classification, with a focus on the impact of SMOTE-Tomek preprocessing. It outlines the error metrics used, compares the baseline performance of the models to their performance with SMOTE-Tomek integration, highlights the interpretability of the logistic regression model, and discusses the achieved improvements and their implications.
III-A Error Metrics
Let TP, FP, TN, and FN denote the true positives, false positives, true negatives, and false negatives, respectively. These metrics serve as the foundational components for evaluating the performance of classification models.
Precision quantifies the proportion of correctly predicted positive cases out of all predicted positive cases:
$\text{Precision}=\frac{TP}{TP+FP},$ (3)
providing an indication of the model’s reliability in identifying true positive instances among its positive predictions.
Recall, also referred to as sensitivity, evaluates the proportion of actual positive cases that the model correctly identifies:
$\text{Recall}=\frac{TP}{TP+FN},$ (4)
measuring the model’s effectiveness in capturing all relevant positive instances.
F1-Score integrates precision and recall into a single performance metric by calculating their harmonic mean:
[…]
offering a balanced evaluation of the model’s precision and recall.
Accuracy assesses the overall correctness of the model by accounting for both positive and negative classifications:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$ (6)
providing a general measure of the model’s performance across all classes.
Matthews Correlation Coefficient (MCC) offers a comprehensive evaluation of binary classification quality, particularly under imbalanced data conditions. It accounts for all elements of the confusion matrix (TP, TN, FP, FN) and is defined as:
$\text{MCC}=\frac{(TP\cdot TN)-(FP\cdot FN)}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$ (7)
where the MCC value ranges from $-1$ to $+1$. 1 Represents perfect classification, 0 indicates random or uninformed classification, and $-1$ Denotes complete misclassification.
III-B Baseline Performance Without SMOTE-Tomek
The experimental results, detailed in Table IV and illustrated in Figure 3, reveal that the Linear SVM model consistently surpassed all other classifiers across the evaluated metrics. It achieved the highest performance, with a mean accuracy of 71.10%, precision of 68.25%, recall of 71.10%, F1-score of 66.47%, and an MCC of 0.5977.
Among the KNN models, configurations with k=3 and k=5 exhibited robust performance, with mean accuracies of 67.69% and 68.42%, respectively, and MCC values closely aligned with the Linear SVM, indicating their effectiveness on this dataset. The Gradient Boosting classifier also demonstrated competitive results, with a mean accuracy of 66.98% and an MCC of 0.5344, indicating its ability to model complex decision boundaries. Conversely, ensemble methods such as AdaBoost and Random Forest exhibited suboptimal performance, with AdaBoost showing particularly low precision of 28.21% and MCC of 0.1963, reflecting limitations in handling class imbalances. Traditional classifiers such as LR and DTs achieved moderate performance levels but were consistently outperformed by more advanced approaches like SVM and KNN.
TABLE IV: Performance For Each Model Without SMOTETomek
Model Mean Accuracy (%) Mean Precision (%) Mean Recall (%) Mean F1-Score (%) Mean MCC
Decision Tree 59.55 $\pm$ 2.72 54.71 $\pm$ 4.08 59.55 $\pm$ 2.72 53.48 $\pm$ 3.15 0.4116 $\pm$ 0.0473
Random Forest 48.71 $\pm$ 1.48 38.26 $\pm$ 8.61 48.71 $\pm$ 1.48 34.11 $\pm$ 2.53 0.1796 $\pm$ 0.0672
SVM (Linear) 71.10 $\pm$ 3.09 68.25 $\pm$ 4.18 71.10 $\pm$ 3.09 66.47 $\pm$ 3.73 0.5977 $\pm$ 0.0456
SVM (RBF) 58.31 $\pm$ 2.05 58.05 $\pm$ 5.33 58.31 $\pm$ 2.05 49.74 $\pm$ 2.88 0.4010 $\pm$ 0.0382
Naive Bayes 51.49 $\pm$ 2.04 45.57 $\pm$ 3.96 51.49 $\pm$ 2.04 38.80 $\pm$ 2.91 0.2566 $\pm$ 0.0510
Logistic Regression 59.85 $\pm$ 2.52 56.85 $\pm$ 5.00 59.85 $\pm$ 2.52 51.85 $\pm$ 3.65 0.4181 $\pm$ 0.0470
KNN (k=3) 67.69 $\pm$ 3.84 66.11 $\pm$ 4.98 67.69 $\pm$ 3.84 64.34 $\pm$ 4.55 0.5478 $\pm$ 0.0586
KNN (k=5) 68.42 $\pm$ 2.69 66.37 $\pm$ 3.29 68.42 $\pm$ 2.69 64.95 $\pm$ 3.12 0.5564 $\pm$ 0.0427
KNN (k=7) 67.80 $\pm$ 2.93 65.07 $\pm$ 3.98 67.80 $\pm$ 2.93 63.84 $\pm$ 3.57 0.5468 $\pm$ 0.0452
Gradient Boosting 66.98 $\pm$ 2.92 66.37 $\pm$ 5.05 66.98 $\pm$ 2.92 63.13 $\pm$ 3.05 0.5344 $\pm$ 0.0446
AdaBoost 49.02 $\pm$ 1.14 28.21 $\pm$ 0.94 49.02 $\pm$ 1.14 33.71 $\pm$ 1.33 0.1963 $\pm$ 0.0472
CatBoost 58.82 $\pm$ 3.11 55.21 $\pm$ 5.54 58.82 $\pm$ 3.11 50.73 $\pm$ 4.32 0.3978 $\pm$ 0.0636
Figure 3: Cross validation performance of ML methods without SMOTE-Tomek.
III-C Enhanced Performance Using SMOTE-Tomek
Following the application of the SMOTE-Tomek preprocessing technique, the LR model demonstrated significant performance improvements, as shown in Figure 4 and Table V. The model achieved a mean accuracy of 76.16% ± 2.58%, a mean precision of 75.31% ± 4.30%, a mean recall of 76.16% ± 2.58%, a mean F1-score of 74.34% ± 2.93%, and a MCC of 0.67 ± 0.03. Performance was further enhanced, with accuracy improving by approximately 1%, through grid search optimization using C=10, L2 penalty, and saga solver. Here, C controls the trade-off between model complexity and training accuracy, with C=10 favoring a closer fit to the data. The L2 penalty regularizes model coefficients to prevent overfitting, while the saga solver is an efficient optimization algorithm designed for large datasets [32].
In comparison, Naive Bayes and Gradient Boosting also performed robustly, with mean accuracies exceeding 70% and MCC values of 0.62 and 0.59, respectively. Conversely, AdaBoost showed significant limitations, achieving a mean accuracy of just 16.10% and an MCC of 0.14, indicating its unsuitability for the current classification task.
Figure 4: Cross validation performance of machine learning method with SMOTE-Tomek.
TABLE V: Performance For Each Model With SMOTETomek
Model Mean Accuracy (%) Mean Precision (%) Mean Recall (%) Mean F1-Score (%) Mean MCC
Decision Tree 41.79 ± 12.44 63.89 ± 7.90 41.79 ± 12.44 40.43 ± 9.53 0.3363 ± 0.0780
Random Forest 59.03 ± 4.50 60.82 ± 6.96 59.03 ± 4.50 58.05 ± 5.39 0.4562 ± 0.0686
SVM Linear 74.20 ± 2.83 73.46 ± 4.58 74.20 ± 2.83 71.44 ± 3.26 0.6428 ± 0.0424
SVM RBF 60.27 ± 2.12 61.34 ± 5.40 60.27 ± 2.12 52.68 ± 3.09 0.4358 ± 0.0346
Naive Bayes 70.48 ± 2.28 75.04 ± 2.42 70.48 ± 2.28 71.17 ± 2.06 0.6286 ± 0.0280
Logistic Regression 76.16 ± 2.58 75.31 ± 4.30 76.16 ± 2.58 74.34 ± 2.93 0.6736 ± 0.0370
KNN (3 Neighbors) 57.79 ± 3.65 69.21 ± 3.19 57.79 ± 3.65 59.59 ± 3.78 0.5063 ± 0.0413
KNN (5 Neighbors) 51.60 ± 4.31 68.56 ± 3.76 51.60 ± 4.31 53.00 ± 5.20 0.4596 ± 0.0451
KNN (7 Neighbors) 46.34 ± 4.45 66.64 ± 4.05 46.34 ± 4.45 47.11 ± 5.69 0.4148 ± 0.0463
Gradient Boosting 70.90 ± 2.77 69.08 ± 4.15 70.90 ± 2.77 67.76 ± 3.56 0.5938 ± 0.0419
AdaBoost 16.10 ± 4.42 12.67 ± 6.97 16.10 ± 4.42 9.06 ± 3.07 0.1422 ± 0.0526
CatBoost 68.21 ± 3.32 68.34 ± 4.55 68.21 ± 3.32 66.37 ± 3.83 0.5633 ± 0.0492
III-D Impact of SMOTE-Tomek on Logistic Regression Interpretability
Understanding feature importance is critical in requirements classification tasks, as it provides actionable insights into the linguistic patterns that define requirement categories. Logistic regression (LR), as an interpretable model, offers a transparent mechanism for analyzing these relationships through its coefficients.
The coefficients presented in Table VI (LR with SMOTE-Tomek) and Table VII (LR without SMOTE-Tomek) highlight the influence of key features (words) in predicting requirement types. These coefficients quantify the strength and direction of association between words and their respective types. Positive coefficients indicate strong alignment with a particular requirement type, while the magnitude reflects the importance of the word in driving classification decisions. As an iterpretable model, LR benefits considerably from preprocessing techniques that address noise and class imbalance. SMOTE-Tomek enhances model performance, stabilizes, and refines coefficients.
The results reveal critical insights into the predictive influence of words and the effect of preprocessing on model behavior. For example, while the word ”seconds” in the ”Functional” type exhibited an extreme negative coefficient (-16.59) without preprocessing, SMOTE-Tomek balanced the dataset and reduced this coefficient to 6.55, ensuring a more consistent representation. Similarly, new influential words such as ”interface” in the ”Look-and-Feel” type emerged after preprocessing, indicating that SMOTE-Tomek enhances the visibility of class-distinguishing features.
The coefficients illustrate the inherent relationship between words and requirement types. High coefficients for words such as ”24x7” in ”Availability” and ”encrypted” in ”Security” highlight their criticality in these types. These insights can guide the development of keyword-based heuristics and inform domain-specific feature engineering.
Furthermore, Figure 5 illustrates the stabilizing effect of SMOTE-Tomek on absolute coefficient values, reducing variability and noise in the model.
TABLE VI: Top Words and Their Logistic Regression Coefficients
with SMOTE-Tomek
Class Top Words Coefficients
Functional allow, meeting, capability 6.55, 3.91, 3.83
Availability available, time, connected 8.67, 5.23, 4.91
Legal comply, regulations, case 8.32, 4.66, 3.88
Look-and-Feel interface, simulate, colors 5.46, 5.05, 4.37
Maintainability maintenance, new, releases 6.43, 5.41, 4.31
Operability server, database, microsoft 6.18, 5.08, 5.06
Performance seconds, let, performance 10.29, 7.03, 5.34
Scalability years, simultaneous, 000 5.88, 5.85, 5.14
Security access, data, security 7.05, 5.68, 5.19
Usability 90, training, easy 5.88, 5.06, 4.43
Fault Tolerance fault, goes, continue 7.20, 4.34, 3.96
Portability windows, portable, 2003 6.85, 5.97, 5.32
TABLE VII: Top Words and Their Logistic Regression Coefficients
without SMOTE-Tomek
Requirement Top Words Coefficients
Functional capability, allow, seconds 17.81, 17.69, -16.59
Availability 24x7, availability, connected 19.22, 14.67, 12.34
Legal comply, meet, regulations 20.25, 18.39, 15.36
Look-and-Feel simulate, colors, drill 18.84, 16.10, 13.74
Maintainability navigator, budget, season 14.93, 14.17, 10.90
Operability encoded, provides, distribution 14.13, 13.98, 13.36
Performance seconds, let, performance 28.51, 24.79, 18.85
Scalability simultaneous, 000, capable 13.29, 12.56, 11.82
Security encrypted, access, role 17.40, 17.03, 15.04
Usability training, 90, help 15.65, 14.39, 14.38
Fault Tolerance loss, fault, downloaded 21.61, 19.65, 16.55
Portability portable, mobile, major 15.97, 12.39, 12.02
Figure 5: Comparison of top word coefficients with and without SMOTE-Tomek (absolute values).
III-E Discussion
The results achieved in this study, with the highest accuracy reaching 76.16% using LR after applying SMOTE-Tomek, highlight both the potential and the limitations of modern ML models for requirements classification. While it demonstrates improvement over the baseline, it reveals the challenges imposed by data scarcity and inherent class imbalance in the PROMISE dataset. These challenges persist despite preprocessing techniques, emphasizing the need for a larger, more representative dataset to better capture the diversity and nuances of requirement types.
The implementation of stratified K-fold cross-validation was important in ensuring a robust evaluation framework. By preserving the original class distribution in each fold, this approach addressed the risk of bias introduced by imbalanced data. Crucially, the validation folds remained untouched by preprocessing. The methodology highlights the importance of rigorous evaluation protocols in validating the effectiveness of class imbalance mitigation strategies.
The integration of SMOTE-Tomek preprocessing improved model performance. This hybrid approach enhanced the ability of models to learn from underrepresented classes, as evidenced by the significant improvement in LR’s metrics, including a 7% increase in MCC compared to the baseline. LR, in particular, benefited from the clearer class boundaries and reduced noise provided by SMOTE-Tomek, achieving superior interpretability and scalability for classification tasks. This hybrid approach enhanced the ability of models to learn from underrepresented classes, as evidenced by the significant improvement in LR’s metrics, including a 7% increase in MCC compared to the baseline.
Despite the overall improvement, several challenges emerged:
  • Ensemble methods such as AdaBoost and Random Forest underperformed significantly, even with SMOTE-Tomek preprocessing. AdaBoost, in particular, achieved a mean accuracy of only 16.10% and an MCC of 0.14, reflecting its inability to adapt effectively to the imbalanced and noisy characteristics of the dataset.
  • Gradient Boosting and Linear SVM demonstrated competitive results. However, their performance gains were relatively modest compared to LR.
  • The PROMISE dataset, while widely used, remains small and constrained in its coverage of requirements types. The restricted variety of classes and the reliance on textual features alone further constrained model generalization.
IV Conclusions
We demonstrated the potential of integrating SMOTE-Tomek preprocessing with modern ML models to address class imbalance in requirements classification. LR, in particular, exhibited the most substantial improvement, achieving 76.16% ± 2.58% accuracy and a 7% increase in the MCC compared to the baseline. The interpretability of LR allowed for an analysis of feature importance, while SMOTE-Tomek preprocessing stabilized coefficients, reduced noise, and improved class representation, thus enabling more robust and meaningful predictions.
These findings emphasize the practicality and scalability of modern ML models for imbalanced text datasets, particularly in resource-constrained environments.
The methodology is applicable to other domains, such as legal or healthcare document analysis, where similar challenges exist. Future work will explore larger datasets, advanced feature representations, and hybrid approaches combining lightweight models with deep learning techniques to further enhance performance and generalizability.